Find the perimeter and area of a right triangle if the shortest side is 9 in. and the longest side is 41 in. Include correct units with each part of your solution.
Answer by nerdybill(7384)

Find the perimeter and area of a right triangle if the shortest side is 9 in. and the longest side is 41 in. Include correct units with each part of your solution.
.
Longest side is the hypotenuse.
Applying Pythagorean theorem:
Let b = third side
9^2 + b^2 = 41^2
81 + b^2 = 1681
b^2 = 1600
b = 40 inches
.
Perimeter then is:
40+9+41 = 90 inches